OpenStudy (anonymous):

Looking for help on this A spherical water storage container with a diameter of 1051 m is filled with water at a rate of 30 m^3/s. At what rate is the area of the surface of the water changing when the height of the water is 800 m from the bottom of the tank?

OpenStudy (whpalmer4):

Can you form an expression for the surface area of the water as a function of the height of the water?

OpenStudy (anonymous):

I hate this section and honestly don't know where to begin. So no.

OpenStudy (anonymous):

sec, I'll take a stab

OpenStudy (whpalmer4):

well, what's the formula for the (surface) area of a circle?

OpenStudy (anonymous):

4pir^2

OpenStudy (whpalmer4):

Right. Now the radius is going to be a function of the height of the water, right? At this point, a good picture is probably going to be useful.

OpenStudy (whpalmer4):

Actually, no 4 in there, that would be the surface area of a sphere, we just want the area of the circle which forms the surface of the water. That's going to be \(A = \pi r^2\), right?

OpenStudy (anonymous):

right

OpenStudy (whpalmer4):

 Cross-sectional view of the tank

OpenStudy (anonymous):

\[\pi r^{2}\]\[\frac{ dh }{ dt }\]

OpenStudy (anonymous):

\[\pi(525.5)^{2}(30)\]

OpenStudy (whpalmer4):

my diagram is a bit misleadingly labeled — I don't think the radius of the surface area is 1051 when the height of the water is 800

OpenStudy (whpalmer4):

that 1051 is referring to the radius of the tank, with the { drawn in

OpenStudy (whpalmer4):

If we look at the sphere with the bottom point as (0,0), 

OpenStudy (whpalmer4):

ack, that's not right. formula should be \(x^2 + (y-1051)^2 = 1051^2\) because the center of that circle is at (0,1051)